# Disambiguation in Conversational Question Answering in the Era of LLM: A Survey

### v1

@@ 5 Open Problems and Challenges
How To Orchestrate? This is one of the research questions we posed at the beginning. To first see why we need to ochestrate among the disambiguation policies, in this paper, we systematically analyze the pros and cons of each disambiguation policy, making us the first to do so to the best of our knowledge. We show the list in Table 5, which are: 1) Automatic: Both QR and LFAG are automatic and do not require human validation, unlike clarifying questions. 2) Additonal LLM Call: For CQA, at least one LLM call is needed for answer generation, and so LFAG could be integrated into that same LLM call. But both QR and ACQ require dedicated LLMs. 3) Visible to User: Rewritten queries are not typically visible to the user, whereas users might notice long-form answers and are definitely aware of clarifying questions. 4) High Risk: Each policy affects different processing steps; for example, QR impacts downstream tasks significantly, as incorrect assumptions can lead to wrong answers. 5) UX Disrupting: Repeated QR does not affect user experience as it is not visible, but too many clarifying questions can vex users. LFAG falls in between, as overly long answers are sometimes unwelcome. As can be seen, each approach has unique strengths and weaknesses, necessitating the need of coordination. The challenge lies in determining when to use which policy. For example, always asking clarifying questions can disrupt UX while always rewriting queries can lead to errors Tanjim et al. (2025).
Opportunities. One promising direction for improving disambiguation systems is the development of an agentic framework Wu et al. (2023); Zeng et al. (2023) that can intelligently select and coordinate various disambiguation strategies. However, this approach faces two key challenges: reliably detecting ambiguities and selecting appropriate disambiguation policies. To address ambiguity detection, recent advances in LLM-based data synthesis and model training can be leveraged Zhang et al. (2024a, b). For policy selection, further enhancement can be achieved through a dedicated reasoning component. Inspired by DeepSeek’s R1 framework Guo et al. (2025), we can simulate various conversation scenarios using different disambiguation policies, calculate rewards based on performance metrics (e.g., answer accuracy, user satisfaction, IR performance), and learn optimal policy selection strategies through iterative refinement. This two-pronged approach—combining specialized ambiguity detection with intelligent policy selection—could lead to a more robust and effective “Disambiguation Agent."
@@ 6 Conclusion
In this paper, we have provided a comprehensive analysis of ambiguity and disambiguation in LLM-based CQA systems through three fundamental research questions. First, we have explored different types of ambiguity and proposed a unified taxonomy using three categories. We also highlighted the challenges of accurately detecting ambiguity, even with LLMs. Next, we have categorized various LLM-based disambiguation approaches and reviewed key benchmark datasets and metrics. Finally, we discussed open challenges and opportunities for LLM-based ambiguity detection and disambiguation strategies. By offering a comprehensive review of current research on ambiguities and disambiguation with LLMs, we hope our survey will contribute to the development of more robust and reliable LLM-based applications.

### v2

@@ 5.2 How To Orchestrate?
This is one of the research questions we posed at the beginning. To first see why we need to ochestrate among the disambiguation policies, in this paper, we systematically analyze the pros and cons of each disambiguation policy, making us the first to do so to the best of our knowledge. We show the list in Table˜5, which are: 1) Automatic: Both QR and LFAG are automatic and do not require human validation, unlike clarifying questions. 2) Additonal LLM Call: For CQA, at least one LLM call is needed for answer generation, and so LFAG could be integrated into that same LLM call. But both QR and ACQ require dedicated LLMs. 3) Visible to User: Rewritten queries are not typically visible to the user, whereas users might notice long-form answers and are definitely aware of clarifying questions. 4) High Risk: Each policy affects different processing steps; for example, QR impacts downstream tasks significantly, as incorrect assumptions can lead to wrong answers. 5) UX Disrupting: Repeated QR does not affect user experience as it is not visible, but too many clarifying questions can vex users. LFAG falls in between, as overly long answers are sometimes unwelcome. As can be seen, each approach has unique strengths and weaknesses, necessitating the need of coordination. The challenge lies in determining when to use which policy. For example, always asking clarifying questions can disrupt UX while always rewriting queries can lead to errors Tanjim et al. (2025). These lead to multiple opportunities which we lay out below.
5.3 Opportunities
The next wave of disambiguation in Conversational QA is being shaped by three emerging trends: agentic orchestration, simulation- and reward-driven policy optimization, and evaluation with LLM-as-a-judge.
Agentic orchestration for disambiguation. Recent generations of LLMs (Meta AI, 2024; OpenAI, 2025b, a) offer longer context windows, more reliable tool use, and reasoning-centric architectures. These advances make it increasingly feasible to deploy multi-agent CQA systems (Dibia et al., 2024; Fourney et al., 2024). Within such frameworks, disambiguation becomes a first-class capability for every agent, but especially for the orchestrator or coordinator agent, which explicitly handles user requests and is responsible for carrying forward the task with the help of other agents.
To operationalize this capability, the orchestrator can leverage agent cards or specifications in tandem with other agents to agree on common nomenclature or establish a constitution. By standardizing specification languages, these can systematically encode explicit policies for query handling and clarification, with Agents.md offering a promising step toward such standardization. For instance, an agent card may include a disambiguation extension that determines when to issue clarification prompts (e.g, which tools to use). Moreover, agentic orchestration can incorporate auxiliary mechanisms, such as memory and verification, to safeguard trajectory alignment. Memory management preserves relevant conversational history (Anthropic, 2025) while filtering out noise, thereby reducing the risk of context drift. LLM verification (discussed separately later) serves as an additional checkpoint, validating whether the system’s chosen path aligns with the intended query resolution.
Conversational engines can further enhance this process by engaging users in real-time clarification. Inspired by Bayesian Experimental Design (BED) (Rainforth et al., 2024), a promising direction in this space, as explored in Kobalczyk et al. (2025), is to actively select questions that maximize expected information gain, shifting from implicit reasoning about the best question to explicit evaluation via sampling from the solution space. Agents with such advanced meta-cognitive skills could eventually infer the most informative questions autonomously, combining static policies from agent cards with dynamic dialogue to robustly handle ambiguous input. Thus, embedding these strategies directly into the orchestrator’s as well as other agents’ layer ensures that disambiguation is not an afterthought, but a modular and transparent component of multi-agent coordination.
Simulation- and reward-driven policy selection. A major opportunity lies in training multiple disambiguation policies shown in Table˜5 using simulation. By generating large-scale ambiguous dialogues and optimizing reward-driven objectives, policies can be tuned not only for task accuracy but also for groundedness, efficiency, and user experience. Inspired by advances in reinforcement learning for reasoning using simulaiton (Guo et al., 2025), such systems can test alternative strategies (clarify vs. rewrite vs. direct answer) and optimize routing controllers accordingly. Mukherjee et al. (2025) push a more practical path forward: rather than relying on SFT or preference-based tuning — both burdened with extra hyper-parameters and indirect reward alignment — their approach shows that QA agents can cut straight to the goal with reward-weighted supervised fine-tuning. This new offline RL objective offers a practical step for fine-tuning an orchestrator to select disambiguation policies more effectively than specification or search-based approach outlined above. Multi-agent RL further enables coordination between other agents with multiple roles (planner, retriever, checker) via shared objectives (Chen et al., 2025).
LLM-as-a-judge for user-centric evaluation. Conventional evaluation metrics such as BLEU, ROUGE, and METEOR are not aligned with the goals of disambiguation, where success requires semantic correctness, underspecification resolution, and conversational coherence. Recent work demonstrates that LLM-as-a-judge can evaluate responses more holistically according to faithfulness, clarity, relevance, and conversational satisfaction (Zheng et al., 2023; Gu et al., 2024; Lee et al., 2025). Specialized judges can enable multilingual and domain-targeted assessment (Kim et al., 2023b; Pombal et al., 2025). Embedding such judges directly “in the loop” (i.e., LLM-verification mentioned earlier) can provide dense, rubric-based feedback not only for outputs but also for policies—e.g., “should a clarification have been asked here?” This can align evaluation goals with end-user satisfaction and accelerate policy refinement. Beyond scoring final answers using traditional LLM-as-a-judge, agent-as-a-judge is another promising direction, where evaluating agents can audit intermediate steps (e.g., query reformulations, retrieval choices, clarification turns) (Zhuge et al., 2025).
Human factors and user experience. User experience remains central to all disambiguation strategies. Adaptive clarification thresholds that are confidence- and risk-aware, persona-sensitive clarification styles, and transparent attribution mechanisms can directly improve user trust and satisfaction. Importantly, minimizing unnecessary interruptions while ensuring correctness is crucial. As LLMs continue to scale in reasoning, planning, and orchestration capabilities, we anticipate that agentic CQA systems will increasingly arbitrate between clarification, rewriting, retrieval, and direct answering and advance toward trustworthy, engaging, reliable multi-agent systems.
@@ 6 Conclusion
In this paper, we have provided a comprehensive analysis of ambiguity and disambiguation in LLM-based CQA systems through three fundamental research questions. First, we have explored different types of ambiguity and proposed a unified taxonomy using three categories. We also highlighted the challenges of accurately detecting ambiguity, even with LLMs. Next, we have categorized various LLM-based disambiguation approaches and reviewed key benchmark datasets and metrics. Finally, we discussed open challenges and opportunities for LLM-based ambiguity detection and disambiguation strategies, particularly from agentic perspectives. By offering a comprehensive review of current research on ambiguities and disambiguation with LLMs, we hope our survey will contribute to the development of more robust and reliable LLM-based applications.
@@
Acknowledgements
This work was supported by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (RS-2023-00216011), as well as another IITP grant funded by the Korea government(MSIT) (RS-2022-II220077).